OpenStudy (amberlynnmaxwell):

Please help :) 23=y+6, what does y equal? Somehow I got 17, but I'm pretty sure that's not correct. Can someone help me figure this out? Thanks to the people that help me!

satellite73 (satellite73):

it is correct

MsBrains (ms-brains):

y=17. You are correct. ^^

MsBrains (ms-brains):

Let's solve your equation step-by-step. 23=y+6 Step 1: Flip the equation. y+6=23 Step 2: Subtract 6 from both sides. y+6−6=23−6 y=17 Answer: y=17

pooja195 (pooja195):

You're correct :-) 23=y+6 to isolate the y you would simply subtract 6 from both sides 23-6=y+6 -6 and you get: y=17

jimthompson5910 (jim_thompson5910):

Here's how to check 23=y+6 23=17+6 ... replace every copy of 'y' with '17' 23 = 23 ... Combine like terms on the right side. The equation is true since the last equation is true, this means the first equation is true when y = 17. So the answer is confirmed.

OpenStudy (amberlynnmaxwell):

Thank you everyone :) I hope you all have a wonderful day ;D

MsBrains (ms-brains):

What I wrote above is not my work btw. It's from an online calculator. Mathpapa .com

OpenStudy (amberlynnmaxwell):

@Ms-Brains that's fine, you still helped me, which I greatly appreciate:)